OpenStudy (anonymous):

Given the equation of the circle (x – 9)2 + y2 = 484, the center of the circle is located at __________, and its radius has a length of __________ units.

OpenStudy (anonymous):

A. (0, 9) and 22 B. (9, 0) and 22 C. (0, 9) and 23 D. (9, 0) and 23

OpenStudy (whpalmer4):

Formula for a circle with radius \(r\) and center \((h,k)\) is \[(x-h)^2+(y-k)^2=r^2\] Read off the values and choose the right answer.

OpenStudy (anonymous):

The center is (9,0) and the radius is : \[\sqrt{484}=22\]